OpenStudy (anonymous):

Help! What is the least common denominator of 4/x^2y+5/y^2?

OpenStudy (jinx.exe):

4/x2y-5/xy2 Final result : -y • (5xy - 4) —————————————— x2 Step by step solution : Step 1 : Raise y to the 2nd power Exponentiaion : Equation at the end of step 1 : 4 5 (————•y)-(—•y2) (x2) x Step 2 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 2 : 4 5y2 (—— • y) - ——— x2 x Step 3 : 4y 5y2 Simplify —— - ——— x2 x Calculating the Least Common Multiple : 3.1 Find the Least Common Multiple The left denominator is : x2 The right denominator is : x Number of times each Algebraic Factor appears in the factorization of: Algebraic Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} x 2 1 2 Least Common Multiple: x2 Calculating Multipliers : 3.2 Calculate multipliers for the two fractions Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = 1 Right_M = L.C.M / R_Deno = x Making Equivalent Fractions : 3.3 Rewrite the two fractions into equivalent fractions Two fractions are called equivalent if they have the same numeric value. For example : 1/2 and 2/4 are equivalent, y/(y+1)2 and (y2+y)/(y+1)3 are equivalent as well. To calculate equivalent fraction , multiply the Numerator of each fraction, by its respective Multiplier. L. Mult. • L. Num. 4y —————————————————— = —— L.C.M x2 R. Mult. • R. Num. 5y2 • x —————————————————— = ——————— L.C.M x2 Adding fractions that have a common denominator : 3.4 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: 4y - (5y2 • x) 4y - 5xy2 —————————————— = ————————— x2 x2 Pulling out like terms : 3.5 Pull out like factors : 4y - 5xy2 = -y • (5xy - 4) Final result : -y • (5xy - 4) —————————————— x2

OpenStudy (anonymous):

Thank you. Could you help me with one more? A/5+a/x+5?

OpenStudy (jinx.exe):

a/5+a/x+5 Final result : ax + 5a + 25x ————————————— 5x Step by step solution : Step 1 : a a Simplify — + — 5 x Calculating the Least Common Multiple : 1.1 Find the Least Common Multiple The left denominator is : 5 The right denominator is : x Number of times each prime factor appears in the factorization of: Prime Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} 5 1 0 1 Product of all Prime Factors 5 1 5 Number of times each Algebraic Factor appears in the factorization of: Algebraic Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} x 0 1 1 Least Common Multiple: 5x Calculating Multipliers : 1.2 Calculate multipliers for the two fractions Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = x Right_M = L.C.M / R_Deno = 5 Making Equivalent Fractions : 1.3 Rewrite the two fractions into equivalent fractions Two fractions are called equivalent if they have the same numeric value. For example : 1/2 and 2/4 are equivalent, y/(y+1)2 and (y2+y)/(y+1)3 are equivalent as well. To calculate equivalent fraction , multiply the Numerator of each fraction, by its respective Multiplier. L. Mult. • L. Num. a • x —————————————————— = ————— L.C.M 5x R. Mult. • R. Num. a • 5 —————————————————— = ————— L.C.M 5x Adding fractions that have a common denominator : 1.4 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: a • x + a • 5 ax + 5a ————————————— = ——————— 5x 5x Equation at the end of step 1 : (ax + 5a) ————————— + 5 5x Step 2 : ax+5a Simplify ————— + 5 5x Rewriting the whole as an Equivalent Fraction : 2.1 Adding a whole to a fraction Rewrite the whole as a fraction using 5x as the denominator : 5 5 • 5x 5 = — = —————— 1 5x Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Pulling out like terms : 2.2 Pull out like factors : ax + 5a = a • (x + 5) Adding fractions that have a common denominator : 2.3 Adding up the two equivalent fractions a • (x+5) + 5 • 5x ax + 5a + 25x —————————————————— = ————————————— 5x 5x Final result : ax + 5a + 25x ————————————— 5x can i get a medal :3 for being fabulous ~